OpenStudy (anonymous):

can someone help me? for the data in the table, does y vary directly with x? if it does, write an equation for the direct variation X Y 8 11 16 22 24 33

OpenStudy (espex):

The first part of this question asks, does x vary with y...so answer that first, how does x change with each change in y?

OpenStudy (anonymous):

3 6 8 ? I don't know im confused

OpenStudy (espex):

You're close, check the math on that last one and I think you'll see that it's not 8.

OpenStudy (anonymous):

ok is it 9?

OpenStudy (espex):

Yes, now if you were to set it up as a table like this, can you see the pattern? If X=8, then y=x+3 if X=16, then y=x+6 if X=24, then y=x+9

OpenStudy (anonymous):

yes I see the pattern

OpenStudy (espex):

Good, so how could you turn that into an equation?

OpenStudy (anonymous):

I don't know y=mx+3 I don't know im still kind of confused

OpenStudy (espex):

That's a good first guess. But look and see that 'x' never changes, you in fact increment the number (n) added to 'x'. So what if you wrote it out like this instead: If X=8, then y=x+(1*3) if X=16, then y=x+(2*3) if X=24, then y=x+(3*3)

OpenStudy (espex):

So what changes, what variable could you use to count up from your 0?

OpenStudy (anonymous):

im sorry but im still not understanding

OpenStudy (espex):

Okay, so you know that you have 'x', and whatever 'x' is equal to, you add to it to get 'y'.

OpenStudy (espex):

You also can see that you are counting up by 1 each time and multiplying that by 3, and then adding it to 'x'.

OpenStudy (espex):

You are counting, 1, 2, 3, 4, .... , all the way to 'n'. So could you write it like this? y = x + 3n, where n is equal to (1, 2, 3, 4, ... n)

OpenStudy (anonymous):

oh I didn't know that's what you were asking yeah I understand that now

OpenStudy (espex):

Yeah, you were very close, I knew you could get it.

OpenStudy (anonymous):

thank you:) but are we done with the problem them?

OpenStudy (espex):

Yes, you found the way that 'y' varied with 'x', and you have created an equation to describe that behavior, y = x + 3n, n=1,2,3,...n

OpenStudy (anonymous):

oh ok now I really understand thank you for your time

OpenStudy (espex):

You're quite welcome. :) Good luck.